OpenStudy (anonymous):

expand this in binomial theorem: (x+1)^5

OpenStudy (anonymous):

\[(x+a)^{n}=\sum_{k=0}^{n}\left(\begin{matrix}n \\ k\end{matrix}\right)x^{k}a^{n-k}\] n=5, a=1

OpenStudy (anonymous):

So the first term, k=0, will be: \[a^{5}\] The second, k=1: \[5xa^{4}\]

OpenStudy (anonymous):

You continue like this until k=5.

OpenStudy (anonymous):

x^5+5x^4+10x^3+10x^2+5x+1